OpenStudy (anonymous):

(5x-1)/2+(-3x-1)/4=9/2 please help i would like to know how to do this problem

OpenStudy (agreene):

(5x-1)/2+(-3x-1)/4=9/2 I prefer: 1/2(5x-1)+1/4(-3x-1)=9/2 now we need to combine everything on the left side... im pretty bad with fractions, so I might have messed it up a bit, lol 7x/4-3/4=9/2 add 3/4 to both sides 7x/4=21/4 divide both sides by 7/4 x=3

OpenStudy (anonymous):

thank you but this is a math website and a website to hit on ppl